Air initially at 100 0F and 100 psia and occupying a volume of 0.5 ft3 undergoes a reversible non-flow constant temperature process such that the final pressure becomes 20 psia. find the work done, ft.lbf.
1
Expert's answer
2020-05-24T16:33:16-0400

The equation of an isothermal process with ideal gas (like air) is


"pV=nRT=const"

The work "W" of a reversible isothermal process is given by


"W=-\int_{V_1}^{V_2}pdV=-nRT\int_{V_1}^{V_2}\frac{dV}{V}=-nRT\ln{\frac{V_2}{V_1}},"

where "V_1, V_2" are the initial and the final volume of the air respectively.

Substitute in the first equation the initial parameters


"nRT=100psia\cdot0.5ft^3=50ft.lbf"

We can find the final volume "V_2" from the first equation


"V_2=\frac{nRT}{p_2},"

"V_2=\frac{50ft.lbf}{20psia}=2.5ft^3."

Substitute to the second equation to find the work done


"W=-50\cdot\ln{\frac{2.5}{0.5}}=-80.5ft.lbf."